Find the value of y in the equation

y=2x^2-x+5 if x =4

thank you

Found 2 solutions by Alan3354, tutorcecilia:
Answer by Alan3354(69443)

Find the value of y in the equation
y=2x^2-x+5 if x =4
---------------
Sub 4 for x
y = 2*4^2 - 4 + 5
y = the sum.
Add those. Make an effort.

Answer by tutorcecilia(2152)

y=2x^2-x+5 if x =4
Plug-in "4" instead of "x" and solve the problem.
.
y=2(4)^2-(4)+5 Simplify using the order of operations (PEMDAS)
y=2(16)-4+5
y=32-4+5
y=28+5
y=33